Part 1 (0.3 point) OSeeHint Oliver, amasked crime fighter, and Malcolm, his villainous counterpart both use arrows on a regular basis. Together, they are the main consumers of arrows in their local market. Their demand schedules for arrows are given below Price per arraw $140 $90 $60 $30 S0 Olivers demand Malcolms demand (arrows per week] arrows per week) 20 40 60 80 30 40 50 60 Using the point plotting tool, interpret the demand schedule to plot fin oints on Oliver and Malcolms combineddemand curve. Then use the straight lins tool to draw the demand curve that connects them, Use a sinele line, not multiple segments. To refer to the graphing tutorial for this question type. please click here Price dol lars per 28 OF 29 QUESTIONS COMPLETED く29,29 VIEW SOLUTION SUBMIT ANSWERMarket demand Part 2 (0.7 point) See Hint What would you expect the market quantity demanded to be at a price of $50? The market demand would be 90 arrows per week. 1st attempt 28 OF 29 QUESTIONS COMPLETED < 29.29 > VIEW SOLUTION SUBMIT ANSWER
Answer #1
price Oliver demand malcom demand

market demand =

oliver +malcom

140 0 0 0+0=0
90 20 30 20+30=50
60 40 40 40+40=80
30 60 50 60+50=110
0 80 60 80+90=140

lHo づ 1 90 o) Co 30 じ 11O g0 L10 )L10 lHo C so,90 J 30 C8o, Lo 30 82o t tO 1리

the above image show the price an markets demand for the good asked in the question with five point in one and line in Other